If a>0 then show that Lin(1/1+na)=0


1
Expert's answer
2021-12-19T14:26:14-0500

Solution:

"\lim_{n\rightarrow \infty} (\dfrac{1}{1+na})
\\=\lim_{n\rightarrow \infty} (\dfrac{1}{1+\infty .a})
\\=\lim_{n\rightarrow \infty} (\dfrac{1}{\infty})
\\=0"

Hence, proved.